# Supersymmetric BMS4 Algebras Revisited: Electric/Magnetic Superalgebras and Free Field Realization

### v1

@@ 3.2 Electric super BMS4 algebra
In Section 2.2, we identified ten electric super BMS4 algebras (2.9s)–(2.9ab). However, in the free field mode construction, only the algebras (2.9s), (2.9t), and (2.9u) admit supersymmetry between spin-0 scalars and spin-$\frac{1}{2}$ fermions. Realizing the other algebras requires introducing higher-spin fermion fields. Since, in the bulk 4D unextended super Poincaré theory, the superpartner of a scalar is a $\frac{1}{2} spin$or, we focus here only on the free field realizations of (2.9s), (2.9t), and (2.9u).
Realization of electric algebra (2.9s)
To realize (2.9s), we introduce bosonic field modes
$\varphi^{(-\frac{3}{4},-\frac{3}{4})},\quad\phi^{(-\frac{1}{4},-\frac{1}{4})},\quad\text{with }[\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{r,\bar{r}},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$ (2.9q)
@@
$\psi^{(-\frac{1}{4},-\frac{3}{4})},\quad\bar{\psi}^{(-\frac{3}{4},-\frac{1}{4})},\quad\text{with }\{\psi^{(-\frac{1}{4},-\frac{3}{4})}_{r,\bar{r}},\bar{\psi}^{(-\frac{3}{4},-\frac{1}{4})}_{s,\bar{s}}\}=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}}.$ (2.9r)
The super BMS4 algebra (2.9s) can then be realized as
$L_{m}=-4\pi^{2}i\sum_{s,\bar{s}}\left(s-\frac{1}{4}m\right)\varphi_{s,\bar{s}}\phi_{m-s,-\bar{s}}-\left(s-\frac{3}{4}m\right)\psi_{s,\bar{s}}\bar{\psi}_{m-s,-\bar{s}},$ (2.9s)
@@
$Q^{(\frac{1}{2},0)}_{r,\bar{r}}=2\sqrt{2}\pi^{2}\sum_{s,\bar{s}}\phi_{s,\bar{s}}\psi_{r-s,\bar{r}-\bar{s}},\quad Q^{(0,\frac{1}{2})}_{r,\bar{r}}=2\sqrt{2}\pi^{2}\sum_{s,\bar{s}}\phi_{s,\bar{s}}\bar{\psi}_{r-s,\bar{r}-\bar{s}}.$ (2.9t)
It is straightforward for the interested reader to check that these constructions satisfy the commutation relations (2.9s) using (2.9q) and (2.9r). The transformations of the field modes under the super BMS4 algebra are summarized in Table 1. For simplicity, the actions of $L$ and $\bar{L}$ are omitted in the table, as the (anti-)holomorphic indices already encode their transformations.
Table 1: Field modes transformation under super BMS4 algebra (2.9s).
@@
$\mathcal{L}=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi-\psi\partial_{0}\bar{\psi}-\bar{\psi}\partial_{0}\psi,$ (2.9y)
where the bosonic part is exactly (2.9n), and $\psi$ and $\bar{\psi}$ are fermionic fields with spin $\pm\frac{1}{2}$ respectively. In terms of Carrollian theories, $\phi$ is also referred as electric scalar, and $\psi$ are referred as electric spinor. The conformal dimensions are $\Delta_{\phi}=\frac{1}{2}$ and $\Delta_{\psi}=\Delta_{\bar{\psi}}=1$. The fermionic fields satisfy the equations of motion
$\partial_{0}\psi=0,\quad\partial_{0}\bar{\psi}=0,$ (2.9z)
@@
$\psi(t,z,\bar{z})=\sum_{r,\bar{r}}\psi_{r,\bar{r}}z^{-r-\frac{3}{4}}\bar{z}^{-\bar{r}-\frac{1}{4}},\quad\bar{\psi}(t,z,\bar{z})=\sum_{r,\bar{r}}\bar{\psi}_{r,\bar{r}}z^{-r-\frac{1}{4}}\bar{z}^{-\bar{r}-\frac{3}{4}}.$ (2.9aa)
@@
$\{\phi(t,z,\bar{z}),\bar{\psi}(t,z^{\prime},\bar{z}^{\prime})\}=i\delta^{2}(z-z^{\prime})$ (2.9ab)
is consistent with the mode commutator (2.9r). In this theory, the bosonic field $\phi$ has one degree of freedom, while $\psi$ and $\bar{\psi}$ together account for one fermionic degree of freedom, as they are canonical conjugates. Hence, the numbers of bosonic and fermionic degrees of freedom are equal. Similarly, in the other constructed theories, the total bosonic and fermionic degrees of freedom match, even if the number of individual fields differs.
Realization of electric algebra (2.9t)
The same bosonic modes (2.9q) and fermionic modes (2.9r) can be used to realize (2.9t). The construction of the BMS4 algebra remains the same as in (2.9s), but the supercharges are now defined differently:
$Q^{(\frac{1}{2},1)}_{r,\bar{r}}=2\sqrt{2}\pi^{2}\sum_{s,\bar{s}}\left(\bar{s}-\frac{3}{4}\bar{r}\right)\phi_{s,\bar{s}}\psi_{r-s,\bar{r}-\bar{s}},\quad Q^{(0,\frac{1}{2})}_{r,\bar{r}}=2\sqrt{2}\pi^{2}\sum_{s,\bar{s}}\phi_{s,\bar{s}}\bar{\psi}_{r-s,\bar{r}-\bar{s}}.$ (2.9ac)
@@
Table 2: Field modes transformation under supercharges of (2.9t).
@@
Realization of electric algebra (2.9u)
The same bosonic and fermionic modes (2.9q) and (2.9r) can also be employed to realize (2.9u), with a distinct construction for the supercharges:
$Q^{(\frac{1}{2},0)}_{r,\bar{r}}=2\sqrt{2}\pi^{2}\sum_{s,\bar{s}}\phi_{s,\bar{s}}\psi_{r-s,\bar{r}-\bar{s}},\quad Q^{(0,\frac{3}{2})}_{r,\bar{r}}=2\sqrt{2}\pi^{2}\sum_{s,\bar{s}}\left(\frac{3}{2}\bar{s}-\frac{3}{4}\bar{r}\right)\phi_{s,\bar{s}}\bar{\psi}_{r-s,\bar{r}-\bar{s}}.$ (2.9ag)
@@
Table 3: Field modes transformation under supercharges of (2.9u).
@@
Interestingly, the theory (2.9y) is sufficiently simple that it accommodates a large number of supercharges. Explicitly counting the conjugation, the field modes in (2.9y) can generate the following supercharges:
$Q^{(\frac{1}{2},0)},\quad Q^{(0,\frac{1}{2})},\quad Q^{(\frac{1}{2},1)},\quad Q^{(1,\frac{1}{2})},\quad Q^{(\frac{3}{2},0)},\quad Q^{(0,\frac{3}{2})},$ (2.9ak)
@@
$\{Q^{(\frac{1}{2},1)}_{r,\bar{r}},Q^{(0,\frac{3}{2})}_{s,\bar{s}}\}=-i(4\pi^{2})\sum_{t,\bar{t}}-\frac{3}{16}(3\bar{r}^{2}+2\bar{r}(\bar{s}-4\bar{t})+2(\bar{s}-2\bar{t})^{2})\phi_{t,\bar{t}}\phi_{r+s-t,\bar{r}+\bar{s}-\bar{t}}.$ (2.9al)
This shows that the anticommutator introduces bosonic generators beyond $M$. Similar structures appear in other anticommutators as well. While these additional generators are indeed symmetries of the specific theory (2.9y), they do not correspond to symmetries of generic BMS4 theories.
@@ 3.3 Magnetic super BMS4 algebra
We now turn to the realization of the magnetic super BMS4 algebras. The Type I algebra (2.9ac) can be constructed using the same modes as in (2.9q) and (2.9r). However, the altered transformation rules under the symmetry imply that these modes correspond to a different physical theory. In contrast, realizing the Type II algebra (2.9ad) requires introducing a distinct set of modes.
Realization of magnetic Type I algebra (2.9ac)
For convenience, we repeat the modes (2.9q) and (2.9r) here, with a renaming of one set of fermionic modes:
$\varphi^{(-\frac{3}{4},-\frac{3}{4})},\quad\phi^{(-\frac{1}{4},-\frac{1}{4})},\quad[\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{r,\bar{r}},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$ (2.9am)
@@
The Type I algebra (2.9ac) is realized as
$L_{m}=-4\pi^{2}i\sum_{s,\bar{s}}\left(s-\frac{1}{4}m\right)\varphi_{s,\bar{s}}\phi_{m-s,-\bar{s}}-\left(s-\frac{3}{4}m\right)\psi_{s,\bar{s}}\chi_{m-s,-\bar{s}},$ (2.9an)
@@
$Q^{(\frac{1}{2},0)}_{r}=-2\sqrt{2}\pi^{2}i\sum_{s,\bar{s}}2\left(\frac{1}{2}s-\frac{1}{4}r\right)\varphi_{s,\bar{s}}\chi_{r-s,-\bar{s}}+\phi_{s,\bar{s}}\psi_{r-s,-\bar{s}},$ (2.9ao)
@@
Notice that the construction of $M$ in (2.9an) differs from that in (2.9s). Under the algebra (2.9ac), the modes transform as summarized in Table 4.
Table 4: Field modes transformation under Type I BMS4 algebra (2.9ac).
@@
$\mathcal{L}=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi-\psi_{+}\partial_{0}\psi_{-}-\psi_{-}\partial_{0}\psi_{+}+\psi_{-}\partial_{z}\psi_{-}.$ (2.9at)
@@
$\partial_{0}^{2}\phi=0,\quad\partial_{0}\psi_{+}=-\partial_{z}\psi_{-},\quad\partial_{0}\psi_{-}=0,$ (2.9au)
@@
Realization of magnetic Type II algebra (2.9ad)
@@
$\varphi^{(-\frac{3}{4},-\frac{3}{4})},\quad\phi^{(-\frac{1}{4},-\frac{1}{4})},\quad[\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{r,\bar{r}},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$ (2.9ax)
@@
The Type II algebra (2.9ad) is realized through the following mode constructions:
$L_{m}=-4\pi^{2}i\sum_{s,\bar{s}}\left(s-\frac{1}{4}m\right)\varphi_{s,\bar{s}}\phi_{m-s,-\bar{s}}-\left(s+\frac{1}{4}m\right)\psi_{s,\bar{s}}\chi_{m-s,-\bar{s}},$ (2.9ay)
@@
$Q^{(\frac{1}{2},0)}_{r}=-4\pi^{2}i\sum_{s,\bar{s}}\varphi_{s,\bar{s}}\chi_{r-s,-\bar{s}}-\left(\frac{1}{2}s-\frac{3}{4}r\right)\phi_{s,\bar{s}}\psi_{r-s,-\bar{s}},$ (2.9az)
@@
Table 5: Field modes transformation under Type II BMS4 algebra (2.9ad).
@@
$\mathcal{L}=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi-\psi_{+}\partial_{0}\psi_{-}-\psi_{-}\partial_{0}\psi_{+}.$ (2.9be)
This theory consists of a single scalar field $\phi$ and two fermionic fields $\psi_{\pm}$ with spins $\pm\frac{1}{2}$. Although the Lagrangian (2.9be) appears similar to the electric super BMS4 field theory (2.9y), it describes a different theory because the conformal dimensions of the fermionic fields differ: $\Delta_{\psi_{+}}=2$ and $\Delta_{\psi_{-}}=0$. The equations of motion read
$\partial_{0}^{2}\phi=0,\quad\partial_{0}\psi_{+}=\partial_{0}\psi_{-}=0,$ (2.9bf)
@@
Realization of magnetic Type I-I algebra (2.9ag)
@@
$\varphi^{(-\frac{3}{4},-\frac{3}{4})},\quad\phi^{(-\frac{1}{4},-\frac{1}{4})},\quad[\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{r,\bar{r}},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$ (2.9bi)
@@
The Type I-I algebra (2.9ag) features an additional bosonic generator $R$. The BMS4 part of the algebra can be realized as
$L_{m}=-4\pi^{2}i\sum_{s,\bar{s}}\left(s-\frac{1}{4}m\right)\varphi_{s,\bar{s}}\phi_{m-s,-\bar{s}}-\left(s-\frac{3}{4}m\right)\psi_{s,\bar{s}}\chi_{m-s,-\bar{s}}$ (2.9bj)
@@
$Q^{(\frac{1}{2},0)}_{r}=-2\pi^{2}i\sum_{s,\bar{s}}2\sqrt{2}\left(\frac{1}{2}s-\frac{1}{4}r\right)\varphi_{s,\bar{s}}\chi_{r-s,-\bar{s}}+\sqrt{2}~\phi_{s,\bar{s}}\psi_{r-s,-\bar{s}}$ (2.9bk)
@@
$R^{(0,0)}_{r,\bar{r}}=\pi^{2}i\sum_{s,\bar{s}}\sqrt{2}~\phi_{s,\bar{s}}\bar{\phi}_{r-s,\bar{r}-\bar{s}}+2~\chi_{s,\bar{s}}\bar{\chi}_{r-s,\bar{r}-\bar{s}}.$ (2.9bl)
The modes transformation is shown in Table 3.3.
Table 6: Field modes transformation under Type I-I BMS4 algebra (2.9ag).
@@
$\mathcal{L}$ $=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi-\psi_{+}\partial_{0}\psi_{-}-\psi_{-}\partial_{0}\psi_{+}+\psi_{-}\partial_{z}\psi_{-}$ (2.9ca)
@@
$\phi:l_{\phi}=0,\quad\Delta_{\phi}=\frac{1}{2},$ (2.9cb)
@@
$\partial_{0}^{2}\phi=0,\quad\partial_{0}\psi_{+}=-\partial_{z}\psi_{-},\quad\partial_{0}\psi_{-}=0,$ (2.9cc)
@@
$\begin{aligned} &\bar{\pi}(t,z,\bar{z})=\sum_{r,\bar{r}}\bar{\varphi}_{r,\bar{r}}z^{-r-\frac{3}{4}}\bar{z}^{-\bar{r}-\frac{3}{4}}-2\left(r+\frac{1}{4}\right)\left(r+\frac{1}{4}\right)\bar{\phi}_{r,\bar{r}}z^{-r-\frac{1}{4}}\bar{z}^{-\bar{r}-\frac{1}{4}},\\ &\bar{\phi}(t,z,\bar{z})=\sum_{r,\bar{r}}\bar{\phi}_{r,\bar{r}}z^{-r-\frac{1}{4}}\bar{z}^{-\bar{r}-\frac{1}{4}}.\end{aligned}$ (2.9cg)
Realization of magnetic Type II-II algebra (2.9ah)

### v2

@@ 3.2 Electric super BMS4 algebra
In Section 2.2, we identified ten electric super BMS4 algebras (2.19)–(2.28). However, in the free field mode construction, only the algebras (2.19), (2.20), and (2.21) admit supersymmetry between spin-0 scalars and spin-$\frac{1}{2}$ fermions. Realizing the other algebras requires introducing higher-spin fermion fields. Since, in the bulk 4D unextended super Poincaré theory, the superpartner of a scalar is a $\frac{1}{2} spin$or, we focus here only on the free field realizations of (2.19), (2.20), and (2.21).
Realization of electric algebra (2.19)
To realize (2.19), we introduce bosonic field modes
$\varphi^{(-\frac{3}{4},-\frac{3}{4})},\quad\phi^{(-\frac{1}{4},-\frac{1}{4})},\quad\text{with }[\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{r,\bar{r}},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$ (3.17)
@@
$\psi^{(-\frac{1}{4},-\frac{3}{4})},\quad\bar{\psi}^{(-\frac{3}{4},-\frac{1}{4})},\quad\text{with }\{\psi^{(-\frac{1}{4},-\frac{3}{4})}_{r,\bar{r}},\bar{\psi}^{(-\frac{3}{4},-\frac{1}{4})}_{s,\bar{s}}\}=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}}.$ (3.18)
The super BMS4 algebra (2.19) can then be realized as
$L_{m}=-4\pi^{2}i\sum_{s,\bar{s}}\left(s-\frac{1}{4}m\right)\varphi_{s,\bar{s}}\phi_{m-s,-\bar{s}}-\left(s-\frac{3}{4}m\right)\psi_{s,\bar{s}}\bar{\psi}_{m-s,-\bar{s}},$ (3.19)
@@
$Q^{(\frac{1}{2},0)}_{r,\bar{r}}=2\sqrt{2}\pi^{2}\sum_{s,\bar{s}}\phi_{s,\bar{s}}\psi_{r-s,\bar{r}-\bar{s}},\quad Q^{(0,\frac{1}{2})}_{r,\bar{r}}=2\sqrt{2}\pi^{2}\sum_{s,\bar{s}}\phi_{s,\bar{s}}\bar{\psi}_{r-s,\bar{r}-\bar{s}}.$ (3.20)
It is straightforward for the interested reader to check that these constructions satisfy the commutation relations (2.19) using (3.17) and (3.18). The transformations of the field modes under the super BMS4 algebra are summarized in Table 1. For simplicity, the actions of $L$ and $\bar{L}$ are omitted in the table, as the (anti-)holomorphic indices already encode their transformations.
Table 1: Field modes transformation under super BMS4 algebra (2.19).
@@
$\mathcal{L}=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi-\psi\partial_{0}\bar{\psi}-\bar{\psi}\partial_{0}\psi,$ (3.21)
where the bosonic part is exactly (3.14), and $\psi$ and $\bar{\psi}$ are fermionic fields with spin $\pm\frac{1}{2}$ respectively. In terms of Carrollian theories, $\phi$ is also referred as electric scalar, and $\psi$ are referred as electric spinor. The conformal dimensions are $\Delta_{\phi}=\frac{1}{2}$ and $\Delta_{\psi}=\Delta_{\bar{\psi}}=1$. The fermionic fields satisfy the equations of motion
$\partial_{0}\psi=0,\quad\partial_{0}\bar{\psi}=0,$ (3.22)
@@
$\psi(t,z,\bar{z})=\sum_{r,\bar{r}}\psi_{r,\bar{r}}z^{-r-\frac{3}{4}}\bar{z}^{-\bar{r}-\frac{1}{4}},\quad\bar{\psi}(t,z,\bar{z})=\sum_{r,\bar{r}}\bar{\psi}_{r,\bar{r}}z^{-r-\frac{1}{4}}\bar{z}^{-\bar{r}-\frac{3}{4}}.$ (3.23)
@@
$\{\phi(t,z,\bar{z}),\bar{\psi}(t,z^{\prime},\bar{z}^{\prime})\}=i\delta^{2}(z-z^{\prime})$ (3.24)
is consistent with the mode commutator (3.18). In this theory, the bosonic field $\phi$ has one degree of freedom, while $\psi$ and $\bar{\psi}$ together account for one fermionic degree of freedom, as they are canonical conjugates. Hence, the numbers of bosonic and fermionic degrees of freedom are equal. Similarly, in the other constructed theories, the total bosonic and fermionic degrees of freedom match, even if the number of individual fields differs.
Realization of electric algebra (2.20)
The same bosonic modes (3.17) and fermionic modes (3.18) can be used to realize (2.20). The construction of the BMS4 algebra remains the same as in (3.19), but the supercharges are now defined differently:
$Q^{(\frac{1}{2},1)}_{r,\bar{r}}=2\sqrt{2}\pi^{2}\sum_{s,\bar{s}}\left(\bar{s}-\frac{3}{4}\bar{r}\right)\phi_{s,\bar{s}}\psi_{r-s,\bar{r}-\bar{s}},\quad Q^{(0,\frac{1}{2})}_{r,\bar{r}}=2\sqrt{2}\pi^{2}\sum_{s,\bar{s}}\phi_{s,\bar{s}}\bar{\psi}_{r-s,\bar{r}-\bar{s}}.$ (3.25)
@@
Table 2: Field modes transformation under supercharges of (2.20).
@@
Realization of electric algebra (2.21)
The same bosonic and fermionic modes (3.17) and (3.18) can also be employed to realize (2.21), with a distinct construction for the supercharges:
$Q^{(\frac{1}{2},0)}_{r,\bar{r}}=2\sqrt{2}\pi^{2}\sum_{s,\bar{s}}\phi_{s,\bar{s}}\psi_{r-s,\bar{r}-\bar{s}},\quad Q^{(0,\frac{3}{2})}_{r,\bar{r}}=2\sqrt{2}\pi^{2}\sum_{s,\bar{s}}\left(\frac{3}{2}\bar{s}-\frac{3}{4}\bar{r}\right)\phi_{s,\bar{s}}\bar{\psi}_{r-s,\bar{r}-\bar{s}}.$ (3.26)
@@
Table 3: Field modes transformation under supercharges of (2.21).
@@
Interestingly, the theory (3.21) is sufficiently simple that it accommodates a large number of supercharges. Explicitly counting the conjugation, the field modes in (3.21) can generate the following supercharges:
$Q^{(\frac{1}{2},0)},\quad Q^{(0,\frac{1}{2})},\quad Q^{(\frac{1}{2},1)},\quad Q^{(1,\frac{1}{2})},\quad Q^{(\frac{3}{2},0)},\quad Q^{(0,\frac{3}{2})},$ (3.27)
@@
$\{Q^{(\frac{1}{2},1)}_{r,\bar{r}},Q^{(0,\frac{3}{2})}_{s,\bar{s}}\}=-i(4\pi^{2})\sum_{t,\bar{t}}-\frac{3}{16}(3\bar{r}^{2}+2\bar{r}(\bar{s}-4\bar{t})+2(\bar{s}-2\bar{t})^{2})\phi_{t,\bar{t}}\phi_{r+s-t,\bar{r}+\bar{s}-\bar{t}}.$ (3.28)
This shows that the anticommutator introduces bosonic generators beyond $M$. Similar structures appear in other anticommutators as well. While these additional generators are indeed symmetries of the specific theory (3.21), they do not correspond to symmetries of generic BMS4 theories.
@@ 3.3 Magnetic super BMS4 algebra
We now turn to the realization of the magnetic super BMS4 algebras. The Type I algebra (2.29) can be constructed using the same modes as in (3.17) and (3.18). However, the altered transformation rules under the symmetry imply that these modes correspond to a different physical theory. In contrast, realizing the Type II algebra (2.30) requires introducing a distinct set of modes.
Realization of magnetic Type I algebra (2.29)
For convenience, we repeat the modes (3.17) and (3.18) here, with a renaming of one set of fermionic modes:
$\varphi^{(-\frac{3}{4},-\frac{3}{4})},\quad\phi^{(-\frac{1}{4},-\frac{1}{4})},\quad[\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{r,\bar{r}},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$ (3.29)
@@
The Type I algebra (2.29) is realized as
$L_{m}=-4\pi^{2}i\sum_{s,\bar{s}}\left(s-\frac{1}{4}m\right)\varphi_{s,\bar{s}}\phi_{m-s,-\bar{s}}-\left(s-\frac{3}{4}m\right)\psi_{s,\bar{s}}\chi_{m-s,-\bar{s}},$ (3.30)
@@
$Q^{(\frac{1}{2},0)}_{r}=-2\sqrt{2}\pi^{2}i\sum_{s,\bar{s}}2\left(\frac{1}{2}s-\frac{1}{4}r\right)\varphi_{s,\bar{s}}\chi_{r-s,-\bar{s}}+\phi_{s,\bar{s}}\psi_{r-s,-\bar{s}},$ (3.31)
@@
Notice that the construction of $M$ in (3.30) differs from that in (3.19). Under the algebra (2.29), the modes transform as summarized in Table 4.
Table 4: Field modes transformation under Type I BMS4 algebra (2.29).
@@
$\mathcal{L}=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi-\psi_{+}\partial_{0}\psi_{-}-\psi_{-}\partial_{0}\psi_{+}+\psi_{-}\partial_{z}\psi_{-}.$ (3.32)
@@
$\partial_{0}^{2}\phi=0,\quad\partial_{0}\psi_{+}=-\partial_{z}\psi_{-},\quad\partial_{0}\psi_{-}=0,$ (3.33)
@@
Realization of magnetic Type II algebra (2.30)
@@
$\varphi^{(-\frac{3}{4},-\frac{3}{4})},\quad\phi^{(-\frac{1}{4},-\frac{1}{4})},\quad[\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{r,\bar{r}},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$ (3.36)
@@
The Type II algebra (2.30) is realized through the following mode constructions:
$L_{m}=-4\pi^{2}i\sum_{s,\bar{s}}\left(s-\frac{1}{4}m\right)\varphi_{s,\bar{s}}\phi_{m-s,-\bar{s}}-\left(s+\frac{1}{4}m\right)\psi_{s,\bar{s}}\chi_{m-s,-\bar{s}},$ (3.37)
@@
$Q^{(\frac{1}{2},0)}_{r}=-4\pi^{2}i\sum_{s,\bar{s}}\varphi_{s,\bar{s}}\chi_{r-s,-\bar{s}}-\left(\frac{1}{2}s-\frac{3}{4}r\right)\phi_{s,\bar{s}}\psi_{r-s,-\bar{s}},$ (3.38)
@@
Table 5: Field modes transformation under Type II BMS4 algebra (2.30).
@@
$\mathcal{L}=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi-\psi_{+}\partial_{0}\psi_{-}-\psi_{-}\partial_{0}\psi_{+}.$ (3.39)
This theory consists of a single scalar field $\phi$ and two fermionic fields $\psi_{\pm}$ with spins $\pm\frac{1}{2}$. Although the Lagrangian (3.39) appears similar to the electric super BMS4 field theory (3.21), it describes a different theory because the conformal dimensions of the fermionic fields differ: $\Delta_{\psi_{+}}=2$ and $\Delta_{\psi_{-}}=0$. The equations of motion read
$\partial_{0}^{2}\phi=0,\quad\partial_{0}\psi_{+}=\partial_{0}\psi_{-}=0,$ (3.40)
@@
Realization of magnetic Type I-I algebra (2.33)
@@
$\varphi^{(-\frac{3}{4},-\frac{3}{4})},\quad\phi^{(-\frac{1}{4},-\frac{1}{4})},\quad[\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{r,\bar{r}},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$ (3.43)
@@
The Type I-I algebra (2.33) features an additional bosonic generator $R$. The BMS4 part of the algebra can be realized as
$L_{m}=-4\pi^{2}i\sum_{s,\bar{s}}\left(s-\frac{1}{4}m\right)\varphi_{s,\bar{s}}\phi_{m-s,-\bar{s}}-\left(s-\frac{3}{4}m\right)\psi_{s,\bar{s}}\chi_{m-s,-\bar{s}}$ (3.44)
@@
$Q^{(\frac{1}{2},0)}_{r}=-2\pi^{2}i\sum_{s,\bar{s}}2\sqrt{2}\left(\frac{1}{2}s-\frac{1}{4}r\right)\varphi_{s,\bar{s}}\chi_{r-s,-\bar{s}}+\sqrt{2}~\phi_{s,\bar{s}}\psi_{r-s,-\bar{s}}$ (3.45)
@@
$R^{(0,0)}_{r,\bar{r}}=\pi^{2}i\sum_{s,\bar{s}}\sqrt{2}~\phi_{s,\bar{s}}\bar{\phi}_{r-s,\bar{r}-\bar{s}}+2~\chi_{s,\bar{s}}\bar{\chi}_{r-s,\bar{r}-\bar{s}}.$ (3.46)
The modes transformation is shown in Table 6.
Table 6: Field modes transformation under Type I-I BMS4 algebra (2.33).
$\varphi_{s,\bar{s}}$ $\phi_{s,\bar{s}}$ $\psi_{s,\bar{s}}$ $\chi_{s,\bar{s}}$
$M_{r,\bar{r}}$ $\phi_{r+s,\bar{r}+\bar{s}}$ 0 $2\left(-\frac{1}{4}r-\frac{1}{2}s\right)\chi_{r+s,\bar{s}}$ 0
$\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}}$ $\frac{1}{\sqrt{2}}\bar{\psi}_{s,\bar{r}+\bar{s}}$ $\sqrt{2}\left(-\frac{1}{4}\bar{r}-\frac{1}{2}\bar{s}\right)\bar{\chi}_{s,\bar{r}+\bar{s}}$ $\bar{\varphi}_{s,\bar{r}+\bar{s}}$ $\left(-\frac{1}{4}\bar{r}-\frac{1}{2}\bar{s}\right)\bar{\phi}_{s,\bar{r}+\bar{s}}$
$\bar{Q}^{(\frac{1}{2},0)}_{r,\bar{r}}$ $\frac{1}{\sqrt{2}}\bar{\chi}_{r+s,\bar{r}+\bar{s}}$ 0 $\left(-\frac{1}{4}r-\frac{1}{2}s\right)\phi_{r+s,\bar{r}+\bar{s}}$ 0
$R^{(0,0)}_{r,\bar{r}}$ $\frac{1}{2\sqrt{2}}\bar{\phi}_{r+s,\bar{r}+\bar{s}}$ 0 $\frac{1}{2}\bar{\chi}_{r+s,\bar{r}+\bar{s}}$ 0
@@
$\mathcal{L}$ $=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi-\psi_{+}\partial_{0}\psi_{-}-\psi_{-}\partial_{0}\psi_{+}+\psi_{-}\partial_{z}\psi_{-}$ (3.47)
@@
$\phi:l_{\phi}=0,\quad\Delta_{\phi}=\frac{1}{2},$ (3.48)
@@
$\partial_{0}^{2}\phi=0,\quad\partial_{0}\psi_{+}=-\partial_{z}\psi_{-},\quad\partial_{0}\psi_{-}=0,$ (3.49)
@@
$\begin{aligned} &\bar{\pi}(t,z,\bar{z})=\sum_{r,\bar{r}}\bar{\varphi}_{r,\bar{r}}z^{-r-\frac{3}{4}}\bar{z}^{-\bar{r}-\frac{3}{4}}-2\left(r+\frac{1}{4}\right)\left(r+\frac{1}{4}\right)\bar{\phi}_{r,\bar{r}}z^{-r-\frac{1}{4}}\bar{z}^{-\bar{r}-\frac{1}{4}},\\ &\bar{\phi}(t,z,\bar{z})=\sum_{r,\bar{r}}\bar{\phi}_{r,\bar{r}}z^{-r-\frac{1}{4}}\bar{z}^{-\bar{r}-\frac{1}{4}}.\end{aligned}$ (3.53)
Realization of magnetic Type II-II algebra (2.34)